OpenStudy (anonymous):

Find the derivative of the function. y = (7x2 + 3) ln(x + 1)

OpenStudy (paxpolaris):

use product rule: \[\left(uv\right)'=u'v+uv'\]

OpenStudy (paxpolaris):

\[y' = 14x \cdot \ln(x+1)+{ 7x^2+3 \over x+1}\]

OpenStudy (anonymous):

ok still not getting it any more help? I need to know if it is -infinity , infinity or -2 , infinity or -infinity, -2 or none of them

OpenStudy (paxpolaris):

is there a limit or ... do you need to evaluate at some value of x ?

OpenStudy (anonymous):

not sure that why I was thinking none of them? It does not give a limit